OpenStudy (anonymous):

14=x+2y and 21= 3x + 10/4y solve for x and y

OpenStudy (luigi0210):

Substitution might work here.

OpenStudy (mathstudent55):

Are these the two equations? Is the second equation correct? \(14 = x + 2y\) \(21 = 3x + \dfrac{10}{4}y\)

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Ok, rewrite the second equation as \(21 = 3x + 2.5y\)

OpenStudy (mathstudent55):

Now multiply the top equation by -3, and add it to the second equation. That will eliminate the x variable.